OpenStudy (anitaasks):

Two trumpet players are playing the same note. If one of them suddenly stops, how much do the sound intensity level decrease by?

OpenStudy (anitaasks):

We are supposed to use the equation, \[\beta(t) = 0dB + 10\log(I/I(0))\] but I have no clue how to get to the answer, which is 3dB

OpenStudy (ookawaiioo):

Decibel (db)=10*log(I/I(0)) Decibel (db)=10*log(2) Log(2) =0.3 10*0.3 = 3 Answer = 3 db or decibels